(a) A 1.05 kg mass is at the center of a uniform spherical shell of mass 1.02 1020 kg and radius 1.08 106 m. Find the resultant gravitational force on the mass.

(b) Find the resultant gravitational force on the mass if it is placed just outside the shell.

magnitude

direction

(c) Find the value of g just outside the shell.
Answer #1

a) inside the spherical shell, gravitational field is zero and hence resultant gravitational force on mass is 0.

b) force outside shell = GMm/R 2

= (6.67×10-11)(1.05)(1.02 × 10 20)/(1.08 × 10 6)2

= 6.12 × 10 -3 N

C) g outside the shell = 6.12/(1.05×1000)

= 5.83 × 10 -3 m/s 2